Your firm is considering a project that would require purchasing $7.6 million worth of new equipment. Determine the present vC and D please
Answer #1

c]

Present value of depreciation tax shield is $2.291 million.

D4 fx =C4/(1+10%) A4 ДА В с D PV of Depreciation Depreciation 1 Year Depreciation tax shield tax shield 0 $1,520,000 $562,400

A B 1 Year 2 0 3 1 42 5 3 64 7 Total Depreciation =7600000*20% =7600000*32% =7600000*19.2% =7600000*11.52% =7600000*11.52% De

d]

Present value of depreciation tax shield = full cost of equipment * tax rate.

Present value of depreciation tax shield = $7.6 million * 35% = $2.660 million.

No discounting is required since the deduction is immediate.